I need help answering questions to the second page.
Problem Set 6A: Production and Costs In the following questions, round to two decimal places (-2). Do not use fractions (4), Rounding Technique: If the third value after the decimal point: Name SOLUTION exceeds "5", round the second value up. Example: 1.456-> 1.46 is less than "5", don't change the second value. Example: 1.454->1.45 5y round the second value to the even value. Eaample. 1455 -146 146-5146 1/2. Given the production function for Tony's Top-Class Tophats (TTT), compute the Marginal Product (MP) and Average Product (AP) values in the table below Labor Total Product Marginal Product .20-효 고 교 교-_a Average Product 2000 22.50 21.62 20.75 1900 17.3316.0 1175 13.%12.50 20 45 65 83 95 104 112 118 122 125 Production/cost information: Each worker costs $54 to hire. In addition, Tony's Top-Class Tophats has total fixed costs of $30 3/7. Complete the following table. Do not use fractions. Round to two decimal places. 3.40 4.233.89 4.56 4.32 Production/cost information: Now suppose costs remain at $30. the cost of each worker rises to $108 to hire. Total fixed 8/10. Complete the following table. Do not use fractions. Round to two decimal places. Cost
Problem Set 6B: Profit Maximization Name: In the following questions, round to two decimal places (2). Use the production and cost information from youe-prevous workeheet. Each worker costs $54 to hire. In addition, Tony's Top-Class Tophats has total fixed costs of $30. 1/2. At a market price of $2.70, the firm will produce and make a total economic PROFIT/ LOSS of $ At a market price of $2.41, the firm will produce units and make a total economic PROFIT/LOSS ofs_ At a market price of $600, the firm will produce units and make a total economic PROFIT/ LOSS of $ If the market price is $6.75 per unit, how many units of output will the firm produce? units 3/4. 5/6. 7. 8. At a market price of $9.00, the firm will make a total economic PROFIT/LOSS of $ 9. The market price is $13.50, and Tony is producing 125 units. To maximize profit he should A. increase output. C. decrease output. B. increase price. D. decrease price. 10. At a market price of $4.00, the firm can make a Problem Set 6C: Profit Maximization Suppose 1/2. At a market price of $1350, the fir n will produce total economic PROFIT/LOSS of that the cost of each worker rises to $108. Total fixed costs remain at $30. units and make a total economic PROFIT / LOSS ofs_ At a market price of $1200, the firm will produce and make a total economic PROFIT / LOSS of $ 3/4. units 5/6. At a market price of $17.00, the firm will produceunits and make a total economic PROFIT/ LOSS of$ If the market price is $6.00 per unit, how many units of output will the firm produce? 7. 8. At a market price of $5.40, the firm will make a 9. The market price is $26.00, and Tony is producing 122 units. To maximize profit he should 10. The market price is $9.00, and the firm is producing 95 units. If total fixed costs decrease by $20, total economic PROFIT/ LOSS of $ A. increase output C. decrease output B. increase price. D. decrease price. Tony shoul INCREASE / DECREASE / NOT CHANGE his output level.
Answer #1

Solution:

Quantity production (or profit maximization) occurs where price equals marginal cost (MC). So, from the tables in Problem Set 6A, we can easily solve for the questions in second image.

With each worker cost of $54 and total fixed costs of $30, we refer to the table 3/7 in problem set 6A. So,

1. At market price of $2.70, the firm will produce at point where marginal cost = $2.70. From table, this results in total product of 20 units or 65 units. The one giving higher profit, or lower loss will be chosen. That occurs when quantity = 65 units.

2. So, total revenue, TR = quantity*price = 65*2.7 = $175.5. Also, at this quantity, total cost, TC = $192

Since, total revenue < total cost, it means that firm will make loss of TC - TR = 192 - 175.5 = $16.5.

3. Similarly,  at market price of $2.41, the firm will produce at point where marginal cost = $2.41. From table, this results in total product of somewhere between 45 and 65 units of output. Let's say output produced = 55 units

4. Total revenue = 2.41*55 = $132.55 and total cost (from table) = somewhere between 138 and 192. So, clearly, loss will occur, though exact magnitude is hard to find.

5. At market price of $6.00, the firm will produce at point where marginal cost = $6.00. From table, this results in total product of 104 units.

6. Total revenue = 6*104 = $624, total cost at that level = $354. So, now since total revenue exceeds total cost, firm makes positive profit of (624 - 354 =) $270

7. At market price of $6.75, total quantity of 112 units will be produced.

8. At market price of $9.00, firm's output will be 118 units, so total revenue = 118*9 = $1,062. Total cost = $462. Clearly, the firm makes huge economic profit.

9. At market price = $13.50, Tony should be producing 122 units to maximize profit, while currently he is producing 125 units. Thus, in order to maximize profit, he should (C) decrease output.